# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v3

@@ Abstract
Using the semiformal method in combinatorics we generalize Pólya’s theorem. This theorem determines the limiting probability that walks in the grid graph $\mathbb{Z}^{d}$ visit the given vertex. We generalize it to the countable complete graph $K_{\mathbb{N}}$ with edge weights in $\mathbb{C}$. In part II we treat edge weights in $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
@@ 1 Introduction
Let $d,n\in\mathbb{N}$ ($\{1,2,\dots\}$) and $\overline{v}\in\mathbb{Z}^{d}$, where $\mathbb{Z}=\{\dots,-1,0,1,\dots\}$ are the integers. Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$ starting at $\overline{0}$ visits in a step $i>0$ the vertex $\overline{v}$. In 1921, G. Pólya proved in [24] a theorem, which we generalize, asserting that for every vertex $\overline{v}$,
@@
Pólya’s theorem is almost always cited in the special case with $\overline{v}=\overline{0}$, although he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a translated quote from [24]). A selection of articles and books discussing Pólya’s theorem is [3, 4, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 29, 30].
@@
$\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\wedge|e|=2\}$}
be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$ on the vertex set $\mathbb{N}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) its cardinality. Let
@@
be a complex edge weight. $\mathbb{C}$ denotes the field of complex numbers, which is endowed with the usual Archimedean absolute value $|z|=\sqrt{z\cdot\overline{z}}$. For $n\in\mathbb{N}_{0}$ let $W(n)$ be the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at the vertex $1$. Thus $W(0)=\{\langle 1\rangle\}$ and for $n>0$ the set $W(n)$ is infinite and countable. For $v\in\mathbb{N}$ let $W(v,n)$ be the subset of $W(n)$ of the walks that visit the vertex $v$ in a step $i>0$. We extend the weight in the usual multiplicative way to
@@
absolutely converges, we denote its sum by $h(W)$. (Results on series are reviewed in the next section.) Suppose that for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges. Then so does $\sum_{w\in W(v,n)}h(w)$ and we consider the generating function
@@
Our generalizations of Pólya’s theorem are formulas for the quantities
$A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,.$}
The formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.24.9 in Section 4. Pólya’s theorem is subsumed in five of them. Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$. Theorems 4.6 (4.7) and 4.9 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
We have fourteen generalizations of one theorem for three reasons. First, in Pólya’s theorem weights are nonnegative real numbers and convergence of infinite series is automatically absolute. For complex weights we consider both kinds of convergence, absolute and conditional one. We also consider, for nonnegative real weights, series with sum $\infty$. Second, Pólya’s theorem lives in the probabilistic setup where weights of edges incident to a fixed vertex sum up to $1$. We consider the case when this condition is kept (convex weights) and the case when it is dropped. Third, already Pólya’s theorem is in reality two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in complexity of proofs. In our generalization we distinguish cases $v=1$ and $v\neq 1$ as well. Now $3\cdot 2\cdot 2=12$; two more generalizations arise by variations of assumptions.
We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we call the set
@@
the image of $C$ by $f$. Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$. We take any bijection $f\colon\mathbb{Z}^{d}\to\mathbb{N}$ such that $f(\overline{0})=1$, and set $v=f(\overline{v})$. The grid graph on $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
@@
determines the weight $h\colon\mathbb{N}_{2}\to\{\frac{1}{2d},0\}$ by the relation $h(e)\neq 0$ iff $e=f[e^{\prime}]$ for some $e^{\prime}\in E_{d}$. Then for every $n\in\mathbb{N}_{0}$ we have
@@
We structure our article as follows. In Section 2 we review infinite series with complex summands and generating functions. Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem. Section 5 contains concluding remarks. We continue our investigation in [17] (see Section 5 for more details).
In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics. It extends the symbolic method in combinatorics [2, 8, 10, 12, 22, 23, 27] from finite to countable sets. In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$. We develop the semiformal method more in [14]. Why “semiformal”? Generating “functions” in the symbolic method are formal power series with coefficients determined by finite operations. Our extension allows operations based on limit transitions. See [15] more applications of the semiformal method.
@@ 2 Series and generating functions in $\mathbb{C}$
We review series and generating functions in the domain $\mathbb{C}$. A series (in $\mathbb{C}$) is a map $h\colon X\to\mathbb{C}$, defined on an at most countable set $X$. We write it as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
@@
    For some $c>0$ and every finite set $Y\subset X$ we have $\sum_{x\in Y}|h(x)|\leq c$.
@@
    If $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$, the limit
    $s=\lim_{n\to\infty}{\textstyle\sum_{i=1}^{n}h(f(i))}$
@@
We call $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$. If the set $X=\{x_{1},x_{2},\dots,x_{n}\}\neq\emptyset$ is finite, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
@@
For $X=\emptyset$ we define the sum as 0. If
$U(x)=\sum_{n\geq 0}u_{n}x^{n}\ \ (\in\mathbb{C}[[x]])$}
is a formal power series, so that $x$ is a formal variable, we denote for any $x\in\mathbb{C}$ by $F_{U}(x)$ the sum of the power series $\sum_{n\geq 0}u_{n}x^{n}$, if it absolutely converges.
@@
such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length $n$ of $w$ is is $n$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
@@
A weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is light if for every $n\in\mathbb{N}_{0}$ the series
@@
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges, but we cannot compare sizes of their sums
@@
is their linear combination. We omit the easy proof of the next result.
@@
If $R$ and $S$ absolutely converge and have respective sums $r$ and $s$, then series $\alpha R+\beta S$ absolutely converges and has the sum $\alpha r+\beta s$.
The next approximation lemma is a key tool.
@@
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$ we have
@@
$|\sum_{n=1}^{N}h(f(n))-s|\leq\frac{\varepsilon}{2}\,\text{ and }\,\sum_{n>N}|h(f(n))|\leq\frac{\varepsilon}{2}$}
and set Y=f[[N]]Y=f[\,[N]\,]. Then for every finite set $Z$ with $Y\subset Z\subset X$ we have
$|s-\sum_{x\in Z}h(x)|\leq|s-\sum_{x\in Y}h(x)|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
@@
Two more operations with series are grouping and product. A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$. If $X$ is at most countable, then so is $P$ and every set $Z\in P$. If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every set $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
@@
Let $A$ and $B$ be sets. We denote their ordered pair by $\langle A,\,B\rangle$. If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are series, then the series
@@
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define more generally their product as the series
@@
where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle$ is an ordered $k$-tuple of $x_{1}$, $\dots$, $x_{k}$ and $\overline{X}=X_{1}\times\dots\times X_{k}$. Iterating Proposition 2.5 we generalize it as follows.
@@
If the series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
@@
If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has the sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for the lightness of the weight $h$ in terms of its values. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk of length at most $n$ and with nonzero weight.
@@
absolutely converges. For $n=0$ it is trivial because $W(0)=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that for every finite set $X\subset W(n-1)$ we have
$\sum_{w\in X}|h(w)|\leq c$}
and that for every $u\in V(h,n-1)$ and every finite set $X\subset\mathbb{N}\setminus\{u\}$ we have
$\sum_{v\in X}|h(\{u,\,v\})|\leq c\,.$}
Let $X\subset W(n)$ be finite. We may assume that walks in $X$ have only edges with nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
@@
Definition 2.10 (convexity)
A light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is convex if for every vertex $u\in V(h)$ the sum
$\sum_{e\in E(h),\,e\ni u}h(e)=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=1\,.$}
We show that these sums exist.
@@
We survey some results on generating functions. We work with univariate formal power series
[…]
They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U(x)$ and $V(x)$ in $\mathbb{C}[[x]]$,
$U(x)+V(x)=\sum_{n\geq 0}(u_{n}+v_{n})x^{n},\ U(x)\cdot V(x)=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}a_{n}x^{n}$ with $a_{0}\neq 0$.
@@
    If the limit
    $U(1)_{1}=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}\ \ (\in\mathbb{C})}$
@@
If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$, but the opposite implication is, in general, not true. For example, if
@@
then $U(1)_{1}$ exists and $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
@@
and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following result is well known.

### v6

@@ Abstract
Using the semiformal method in combinatorics we generalize Pólya’s theorem. This theorem determines limits of probabilities that walks in the grid graph $\mathbb{Z}^{d}$ starting at the origin visit the given vertex. We generalize it to the countable complete graph $K_{\mathbb{N}}$ with edge weights in $\mathbb{C}$. In part II we consider analogs of Pólya’s theorem for edge weights in $\mathbb{C}[[x_{1},\dots,x_{k}]]$.
Besides the abstract, we want to help the reader understand our article by the following preamble. In it we have more space to explain what we do.
We mention Pólya’s theorem, which is a well known result in the discipline of random walks. However, our article is not really about generalizing Pólya’s theorem. We use extensions of Pólya’s theorem to edge weighted countable complete graphs merely as a test case for the novel generalization of the well known ([3, 8, 10, 12, 22, 23, 28]) symbolic method in enumerative (analytic) combinatorics. This generalization is due to the first author, and we call it the semiformal method. It will be developed in detail in [14]. The present article is an explanation of the functioning of the semiformal method in the form of an extensive example. More applications of it are given in [15]. The idea of the semiformal method is to extend finite sums defining coefficients of generating functions in the symbolic method to infinite (countable) series. Finite sums in a ring (in the symbolic method, one also needs to multiply) are straightforward, but it is less clear what is a sum of an infinite series of real or complex numbers. We use three kinds of sums of infinite series: absolute convergence (which is used in probability where summands are nonnegative), the more general and weaker conditional convergence, and sums equal to $\pm\infty$. The last sums are in fact important — the heart of the combinatorial proof of the case $d\leq 2$ in Pólya’s theorem is the fact that a certain series has sum $\infty$. For details on the uses of these three kinds of sums we refer to the article below. But why confine ourselves to infinite series of real or complex numbers, when non-Archimedean infinite series are also interesting? This is the topic of [17]. To summarize, the main contribution of this article is the demonstration of the working of the semiformal method in the fourteen cases that arise when Pólya’s theorem is generalized to $K_{\mathbb{N}}$ with complex edge weights.
@@ 1 Introduction
Let $d,n\in\mathbb{N}$ ($\{1,2,\dots\}$) and $\overline{v}\in\mathbb{Z}^{d}$, where $\mathbb{Z}=\{\dots,-1,0,1,\dots\}$ are the integers. Let $P_{d}(\overline{v},n)$ be the probability that a walk of length $n$ in the grid graph $\mathbb{Z}^{d}$ starting at $\overline{0}$ visits in a step $i>0$ the vertex $\overline{v}$. In 1921, G. Pólya proved in [24] a theorem, which we generalize, that for every vertex $\overline{v}$,
@@
Pólya’s theorem is almost always mentioned in the special case with $\overline{v}=\overline{0}$, but he proved it for any vertex $\overline{v}$ (see [16, Section 1] for a quote where he states his result). A selection of articles and books discussing Pólya’s theorem is [4, 5, 6, 9, 11, 13, 18, 19, 20, 21, 25, 26, 27, 31, 32].
@@
$\mathbb{N}_{2}=\binom{\mathbb{N}}{2}=\{e\colon\;e\subset\mathbb{N}\text{ with }|e|=2\}$}
be edges of the countable complete graph $K_{\mathbb{N}}=\langle\mathbb{N},\mathbb{N}_{2}\rangle$. Here $|A|$ ($\in\mathbb{N}_{0}=\{0,1.\dots\}$) denotes the cardinality of a finite set $A$. Let
@@
be an edge weight, where $\mathbb{C}$ denotes the field of complex numbers, which is endowed with the usual Archimedean absolute value $|z|=\sqrt{z\cdot\overline{z}}$. Let $n\in\mathbb{N}_{0}$. We denote by $W(n)$ the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$. Thus $W(0)=\{\langle 1\rangle\}$ and for $n>0$ the set $W(n)$ is infinite and countable. Let $v\in\mathbb{N}$. We denote by $W(v,n)$ ($\subset W(n)$) the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$ and visiting $v$ in a step $i>0$. We extend the weight multiplicatively to
@@
absolutely converges, we denote its sum by $h(W)$. (Results on series are reviewed in the next section.) Suppose that for every $n\in\mathbb{N}_{0}$ the series $\sum_{w\in W(n)}h(w)$ absolutely converges. Then so does the series $\sum_{w\in W(v,n)}h(w)$ and we consider the generating function
@@
Our generalizations of Pólya’s theorem are formulas for three quantities:
$\lim_{x\nearrow 1}A_{h,v}(x),\ A_{h,v}(1)=\sum_{n\geq 0}h(W(v,\,n))\,\text{ and }\,\lim_{n\to\infty}h(W(v,\,n))\,$,}
where $\lim_{x\nearrow 1}$ means that $x$ goes to $1$ through values $x\in[0,1)$. In some cases, these quantities are twisted by powers $\xi^{n}$ for $\xi\in\mathbb{C}^{*}$ ($\mathbb{C}\setminus\{0\}$). The formulas appear in fourteen theorems: Theorems 3.23.7 in Section 3 and Theorems 4.34.10 in Section 4. Pólya’s theorem is subsumed in five of them. Theorems 3.5 and 3.7 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}=\overline{0}$. Theorems 4.7 (4.8) and 4.10 extend cases $d\geq 3$ and $d\leq 2$, respectively, when $\overline{v}\neq\overline{0}$.
The fourteen generalizations come from three sources. First, weights in Pólya’s theorem are nonnegative real numbers, and the convergence of infinite series is absolute. For our complex weights, we consider two kinds of convergence: absolute and conditional. We also consider real series with sums $\pm\infty$. Second, Pólya’s theorem is on probabilities, which means that the weights of edges incident to a fixed vertex sum up to $\xi=1$. We generalize it to $\xi$-convexity for any $\xi\in\mathbb{C}$; the case of $\xi\neq 0$ is normalized to $|\xi|=1$ and the case $\xi=0$ is treated separately (it corresponds to Kirchhoff’s first law). Thus we consider two cases, general weights and $\xi$-convex weights. Third, Pólya’s theorem is, in reality, two theorems because the cases $\overline{v}=\overline{0}$ and $\overline{v}\neq\overline{0}$ differ in the complexity of proofs. In our generalization, we correspondingly distinguish between cases $v=1$ and $v\neq 1$. Now $3\cdot 2\cdot 2=12$ and two additional generalizations arise by varying the assumptions.
We explain how our approach includes Pólya’s original problem. If $f\colon A\to B$ is a map and $C$ is any set (not necessarily a subset of $A$), we denote by
@@
the image of $C$ by $f$. Let $d\in\mathbb{N}$ and $\overline{v}\in\mathbb{Z}^{d}$. We take any bijection $f\colon\mathbb{Z}^{d}\to\mathbb{N}$ such that $f(\overline{0})=1$, and set $v=f(\overline{v})$. The grid graph $\mathbb{Z}^{d}$, which is the graph $\langle\mathbb{Z}^{d},E_{d}\rangle$ with the edges
@@
determines a weight $h\colon\mathbb{N}_{2}\to\{\frac{1}{2d},0\}$ by the relation $h(e)\neq 0$ iff $e=f[e^{\prime}]$ for $e^{\prime}\in E_{d}$. Then for every $n\in\mathbb{N}_{0}$ we have
@@
See Section 5 for the importance of (the general form of) this equality.
We structure our article as follows. In Section 2 we review infinite series with complex summands and the domain $\mathbb{C}[[x]]$ of generating functions. Sections 3 and 4 contain our main results, the mentioned fourteen generalizations of Pólya’s theorem. Section 5 contains concluding remarks. We continue our investigation in [17]; see Section 5 for more details on [17].
In Sections 3 and 4 we use what we term the semiformal method in (enumerative) combinatorics. It extends the symbolic method in combinatorics [3, 8, 10, 12, 22, 23, 28] from finite to countable sets. In our situation such extension is natural and inevitable, coefficients in generating functions $A_{h,v}(x)$ arise as sums of infinite series $\sum_{w\in W(v,n)}h(w)$. We develop this method more in [14]. Why “semiformal”? Generating “functions” in the symbolic method are formal power series and their coefficients are determined by finite operations. We also use formal power series, but our extension allows operations based on limit transitions. See [15] for more applications of the semiformal method.
@@ 2 Series and generating functions in $\mathbb{C}$
We review series and generating functions in $\mathbb{C}$. A series is a map $h\colon X\to\mathbb{C}$ defined on a finite or countable set $X$; we write it as $\sum_{x\in X}h(x)$. We say that $\sum_{x\in X}h(x)$ absolutely converges if two equivalent conditions hold.
@@
    For some constant $c\geq 0$ and every finite set $Y\subset X$ we have
    $\sum_{x\in Y}|h(x)|\leq c\,.$}
@@
    If the set $X$ is countable, then for every bijection $f\colon\mathbb{N}\to X$ the proper limit
    $s=\lim_{n\to\infty}{\textstyle\sum_{i=1}^{n}h(f(i))\ \ (\in\mathbb{C})}$
@@
(In fact, the independence of $s$ on $f$ may be omitted from the formulation.) We call $s$ the sum of the series and denote it again by $\sum_{x\in X}h(x)$. If the set $X$ is finite and nonempty, $X=\{x_{1},x_{2},\dots,x_{n}\}$, then every series $\sum_{x\in X}h(x)$ absolutely converges and has the sum
@@
If $X=\emptyset$, we define the sum as 0. If
$U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\ \ (\in\mathbb{C}[[x]])$}
is a formal power series, so that $x$ is a formal variable, we denote for any $c\in\mathbb{C}$ by $F_{U}(c)$ the sum of the power series $\sum_{n\geq 0}u_{n}c^{n}$, if it absolutely converges. Let
$H_{1}={\textstyle\{U(x)\in\mathbb{C}[[x]]\colon\;\text{$U(c)$ absolutely converges for every $c\in\mathbb{C}_{<1}$}\}\,,}$
where $\mathbb{C}_{<1}=\{c\in\mathbb{C}\colon\;|c|<1\}$. Thus for any $U\in H_{1}$ the sums form the function $F_{U}\colon\mathbb{C}_{<1}\to\mathbb{C}$.
@@
such that $n\in\mathbb{N}_{0}$ and $v_{i-1}\neq v_{i}$ for every $i\in[n]$ ($\{1,2,\dots,n\}$, $0]=\emptyset$). The length of $w$ is $n$ ($\in\mathbb{N}_{0}$). We extend $h$ to walks. For $n>0$ we set
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$. We call the weight $h$ light if for every $n\in\mathbb{N}_{0}$, the series
@@
Let $S=\sum_{x\in X}h(x)$ be a series. If $Y\subset X$, we say that $\sum_{x\in Y}h(x)$ is a subseries of $S$. Any subseries of an absolutely convergent series absolutely converges. However, we cannot compare the sizes of sums
@@
is the linear combination of $R$ and $S$. We omit the easy proof of the next result.
@@
If $R$ and $S$ absolutely converge and have sums $r$ and $s$, respectively, then $\alpha R+\beta S$ absolutely converges and has sum $\alpha r+\beta s$.
The next approximation lemma is useful.
@@
If $S=\sum_{x\in X}h(x)$ absolutely converges and has sum $s$, then for every $\varepsilon>0$ there exists a finite set $Y\subset X$, denoted by $Y(S,\,\varepsilon)$, such that for every finite set $Z$ with $Y\subset Z\subset X$,
@@
$\big|\sum_{n=1}^{N}h(f(n))-s\big|\leq\frac{\varepsilon}{2}\,\text{ and }\,\sum_{n>N}|h(f(n))|\leq\frac{\varepsilon}{2}\,$,}
and set $Y=\{f(1),f(2),\dots,f(N)\}$. Then for every finite set $Z$ with $Y\subset Z\subset X$ we have
$\big|s-\sum_{x\in Z}h(x)\big|\leq\big|s-\sum_{x\in Y}h(x)\big|+\sum_{x\in Z\setminus Y}|h(x)|\leq\frac{\varepsilon}{2}+\frac{\varepsilon}{2}=\varepsilon\,.$}
@@
Two operations on series are grouping and product. A partition of a set $X$ is a set $P$ of nonempty and disjoint sets such that $\bigcup P=X$. If $X$ is at most countable, then so is $P$ and every set $Z\in P$. If $S=\sum_{x\in X}h(x)$ is a series and $P$ is a partition of $X$ such that for every $Z\in P$ the subseries $\sum_{x\in Z}h(x)$ absolutely converges and has sum $s_{Z}$, then the series
@@
Let $X_{1}$, $X_{2}$ be sets and $X_{1}\times X_{2}$ be their Cartesian product. If $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i=1,2$, are series, then the series
@@
Let $S_{i}=\sum_{x\in X_{i}}h_{i}(x)$, $i\in[k]$, be $k\in\mathbb{N}$ series. We define their product as the series
@@
where $\overline{x}=\langle x_{1},\dots,x_{k}\rangle\in X_{1}\times\dots\times X_{k}=\overline{X}$. Iterating Proposition 2.5 we obtain the following result.
@@
If series $S_{1}$, $\dots$, $S_{k}$ absolutely converge and have sums $s_{1}$, $\dots$, $s_{k}$, respectively, then $\prod_{i=1}^{k}S_{i}$ absolutely converges and has sum $s_{1}s_{2}\dots s_{k}$.
@@
If nonempty series $S_{1}$ and $S_{2}$ are not identically zero and if their product $S_{1}\cdot S_{2}$ absolutely converges and has sum $s$, then $S_{1}$ and $S_{2}$ absolutely converge and their respective sums $s_{1}$ and $s_{2}$ satisfy $s_{1}s_{2}=s$.
@@
Let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ and $n\in\mathbb{N}_{0}$. We obtain a sufficient condition for lightness of the weight $h$ in terms of its values. We denote by $V(h,n)$ the set of vertices in $K_{\mathbb{N}}$ reachable from $1$ by a walk with nonzero weight and length at most $n$.
@@
absolutely converges. For $n=0$ it is trivial because $W(0)=\{\langle 1\rangle\}$. Let $n>0$ and let $c>0$ be a constant such that
$\sum_{w\in X}|h(w)|\leq c\,\text{ and }\,\sum_{v\in X^{\prime}}|h(\{u,\,v\})|\leq c$}
for every finite set $X\subset W(n-1)$, every $u\in V(h,n-1)$ and every finite set $X^{\prime}\subset\mathbb{N}\setminus\{u\}$. Let $X\subset W(n)$ be finite. We may assume that all walks in $X$ have nonzero weights. We decompose every walk $w\in X$ as $w=w^{\prime}\,\ell(w)$, where $w^{\prime}\in W(n-1)$ and $\ell(w)$ is the last vertex. Let $Y$ be the set of walks $w^{\prime}$. For $w^{\prime}\in Y$ let
@@
Definition 2.10 ($\xi$-convexity)
Let $\xi\in\mathbb{C}$. We call a light weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ $\xi$-convex if for every vertex $u\in V(h)$ the sum
$\sum_{e\in E(h),\,e\ni u}h(e)=\sum_{v\in\mathbb{N}\setminus\{u\}}h(\{u,\,v\})=\xi\,.$}
We say that $h$ is convex if it is $1$-convex.
We show that the sums in the definition exist.
@@
We survey results on generating functions. We use univariate formal power series $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}$ with complex coefficients. They form an integral domain denoted by $\mathbb{C}[[x]]$. For $U$ and $V$ in $\mathbb{C}[[x]]$,
$U+V=\sum_{n\geq 0}(u_{n}+v_{n})x^{n}\,\text{ and }\,U\cdot V=\sum_{n\geq 0}\big(\sum_{j=0}^{n}u_{j}v_{n-j}\big)x^{n}\,.$}
Neutral elements are $0=0+0x+0x^{2}+\dots$ and $1=1+0x+0x^{2}+\cdots$. Units in $\mathbb{C}[[x]]$ are exactly the formal power series $\sum_{n\geq 0}u_{n}x^{n}$ with $u_{0}\neq 0$.
Let $\mathbb{C}_{1}=\{x\in\mathbb{C}\colon\;|x|=1\}$ and $\xi\in\mathbb{C}_{1}$. We define
$\xi\mathbb{R}_{\geq 0}=\{\xi x\colon\;x\in\mathbb{R}_{\geq 0}\}\ \ (\subset\mathbb{C})\,.$
It is the ray in $\mathbb{C}$ going from the origin in the direction $\xi$. Clearly, $\xi\mathbb{R}_{\geq 0}$ is closed under addition. If $\alpha_{j}\in\xi_{j}\mathbb{R}_{\geq 0}$, $j=1,2$, then $\alpha_{1}\alpha_{2}\in(\xi_{1}\xi_{2})\mathbb{R}_{\geq 0}$. If
$h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$
is a light weight and $n\in\mathbb{N}_{0}$ then $h(W(n))\in\xi^{n}\mathbb{R}_{\geq 0}$. The lines $\xi\mathbb{R}$ in $\mathbb{C}$ going through the origin in the direction $\xi$ have the same properties. For $\alpha\in\mathbb{C}$ and $U=U(x)=\sum_{n\geq 0}u_{n}x^{n}\in\mathbb{C}[[x]]$ we define
$(U)_{\alpha}=(U)_{\alpha}(x):=U(\alpha x)=\sum_{n\geq 0}u_{n}\alpha^{n}x^{n}\ \ (\in\mathbb{C}[[x]])\,.$}
Note that $(U+V)_{\alpha}=(U)_{\alpha}+(V)_{\alpha}$ and $(U\cdot V)_{\alpha}=(U)_{\alpha}\cdot(V)_{\alpha}$.
@@
    If the (proper) limit
    $U(1)_{1}:=\lim_{n\to\infty}{\textstyle\sum_{j=0}^{n}u_{j}\ \ (\in\mathbb{C})}$
@@
If $U(1)_{2}$ exists, then $U(1)_{1}$ exists and $U(1)_{1}=U(1)_{2}$. The opposite implication is, in general, not true. For example, if
@@
then $U(1)_{1}=\log 2$, but $U(1)_{2}$ does not exist. But if
@@
and $U(1)_{1}$ exists, then $U(1)_{2}$ exists and $U(1)_{1}=U(1)_{2}$. The following is a corollary of the well known Hadamard’s formula for the radius of convergence of a formal power series.